OpenStudy (anonymous):

Alex has 20m barb wire to fence a rectangular lot. What are the dimensions of the lot that will contain the maximum area?

OpenStudy (anonymous):

Are you supposed to use calculus?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):



OpenStudy (anonymous):

\[ 20=2x+2y \implies y=\frac{20-2x}{2} \]

OpenStudy (anonymous):

\[ A = xy = x(10-x) \]

OpenStudy (anonymous):

Minimize the function.

OpenStudy (anonymous):

so sir, the answer is ?

OpenStudy (anonymous):

Hell no.

OpenStudy (anonymous):

I did the algebra, sir. You do the Calculus.

OpenStudy (anonymous):

What is \(dA/dx\)?

OpenStudy (anonymous):

integration ..

OpenStudy (anonymous):

Differentiation.

OpenStudy (anonymous):

differential calculus

OpenStudy (anonymous):

\[ A = x(10-x) \]

OpenStudy (anonymous):

Find \(A'\)

OpenStudy (anonymous):

thank u sir.

OpenStudy (ikram002p):

lol use differential !

OpenStudy (anonymous):

You need \(A'=0\) to find critical points.

OpenStudy (ikram002p):

Based on the analysis of @wio \[y=10-x\] \[xy=10x-x\] \[x(10-x)=10x-x\] \[x^2-x=0 \rightarrow the final function\]

OpenStudy (anonymous):

Well it is \[ 10x-x^2 \]And \[ A'=10-2x = 0 \implies x= 5 \]

OpenStudy (ikram002p):